[Hallc_running] [New Logentry] Hall C RC Update
latif at jlab.org
latif at jlab.org
Fri Mar 15 09:55:06 EDT 2019
- Yesterday kinematic group-37 was completed around 3:00 AM.
- Yesterday around 10 AM we tried to do the BCM calibration. It was canceled/postponed later since we required the slit to be closed for zero currents, and the slit was being shared with Hall D. The timing and coordination to repeat the BCM calibration to be discussed.
- We are taking production data smoothly except for some minor issues which were addressed immediately.
- We are currently taking data for kinematic setting 38-5.
- Once done with kinematic group 38, we will move to the kinematic groups 39 and 40 (updated recently).
- CSV will be over around 3:30 PM on Sunday. In preparation for installation on Monday, we will ramp down all the HMS and SHMS magnets, rotate the HMS to about 40 degrees, rotate the SHMS to about 20 degrees. The target expert (David Meekins) will come in to warm up the target in the afternoon. Radcon and MCC have been informed to support the access.
- Detector and beam pipe installation work will take place starting Sunday swing through at least Tuesday day shift.
- MCC informed us that because of the required changes in the configuration, they would like to do the pass change (3 pass) after the RF recovery on Wednesday.
- Shifts starting from Sunday (March 17th) swing shift through at least Tuesday (March 18th) day shifts are canceled. Shift takers on Tuesday Swing shift are on standby and should check e-mail prior to coming to the counting-house. More shifts are likely to be canceled.
- We will have the RC meeting today at 4:30 PM in the counting house to discuss the update and plan to move forward. Bluejeans Conference ID : 861439920
Today's run plan: https://hallcweb.jlab.org/wiki/index.php/RC_meeting_and_Run_Plan,_Friday_March_15
Online analysis instruction: https://hallcweb.jlab.org/wiki/index.php/CSV_Fall_2018_Run_Plan
Information for Shift Takers: https://hallcweb.jlab.org/wiki/index.php/Information_for_Shift_Takers
We will have the RC meeting today at 4:30 PM in the counting house to discuss the update and plan to move forward. Bluejeans Conference ID : 861439920
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3668173
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running